OpenStudy (anonymous):

I=prt simple interstest find the value of I given that p=400, r= 0.11, and t=2

OpenStudy (anonymous):

@mipie13

OpenStudy (anonymous):

so do i multiply again?

OpenStudy (anonymous):

88

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

opps

OpenStudy (anonymous):

I would multiply the whole numbers first.

OpenStudy (anonymous):

400x2=800 then 800x0.11=88

OpenStudy (anonymous):

l= 400*2*0.11 l=800*0.11

OpenStudy (anonymous):

so its 88

OpenStudy (anonymous):

Oops. Well, we posted the same thing basically.

OpenStudy (anonymous):

My bad

OpenStudy (anonymous):

so 88

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Yup.